OpenStudy (anonymous):

PLEASE HELP PEEPS p(x)=4x^3-3x^2+bx+6 x-1 and x+3 are the factors. determine value of b when remainder is same for both factors!

OpenStudy (anonymous):

the remainder is \[p(1)\] and also \[p(-3)\] find these, set the result equal, and solve for b

OpenStudy (anonymous):

hope it is clear what i mean.

OpenStudy (anonymous):

k so i got 7 and 43

OpenStudy (anonymous):

hold on. your answer must have a b in it because you do not know what b is

OpenStudy (anonymous):

for example, \[p(1)=4-3+b+6=b+7\]

OpenStudy (anonymous):

and \[p(-3)=4\times (-3)^3-3\times (-3)^2-3b+6=-3b-129\] if my arithmetic is correct that means \[b+7=-3b+29\] and you can solve that for b

OpenStudy (anonymous):

oh okokok thankyuverymuch goodsir

OpenStudy (anonymous):

I think the calculation is not correct, however the procedure is perfect \[b = -34\]

OpenStudy (anonymous):

howyu get -34

OpenStudy (anonymous):

typo on last line but again my calculation could be off it should be \[b+7=-3b-129\] \[4b=-136\] \[b=-34\] what alsamixer said

OpenStudy (anonymous):

ok